OpenStudy (anonymous):

find the equation of the line containing the given pair of points. (-5,-2) and (-1,-8) what is the equation? y=

OpenStudy (whpalmer4):

We can find the slope, m, by taking the rise between the two points (difference in y values) over the run (difference in x values). then we use one of the points in the equation y = mx + b to find b. difference in y = (-2 - -8) = 6 difference in x = (-5 - -1) = -4 so m = -6/4 -2 = -6/4(-5) + b -2 = 30/4 + b -8/4 - 30/4 = b = -9.5 y = -1.5x-9.5 check with the other point: -8=-1.5(-1)-9.5